OpenStudy (anonymous):

The wildflowers at a national park have been decreasing in numbers. There were 1200 wildflowers in the first year that the park started tracking them. Since then, there have been one fourth as many new flowers each year. Create the sigma notation showing the infinite growth of the wildflowers and find the sum, if possible.

OpenStudy (anonymous):

year new flowers 1 1200 2 300 3 75

OpenStudy (anonymous):

The problem is divergent right?

OpenStudy (lxelle):

First figure out the info given : a1=1200 r=14

OpenStudy (anonymous):

that its not divergent

OpenStudy (lxelle):

since |r| = 1/4 which is less than 1, the series is Convergent

OpenStudy (anonymous):

can i show you what i think the answer is?

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty}1200(\frac{ 1 }{ 4})^{i-1}; 1600 wildflowers\]

OpenStudy (lxelle):

bingo!

OpenStudy (anonymous):

thanks for your help (:

OpenStudy (lxelle):

No prblm.